OpenStudy (anonymous):

i need help badly!! Solve this system of equations. Write the answer as an ordered pair. -5x - 2y = -2 x - 7y = -7

OpenStudy (anonymous):

what did u come up with...and show ur work

OpenStudy (anonymous):

for some reason i thought it was asking for the equation...all i have is..... -5(7y-7)-2y=-2

OpenStudy (anonymous):

multiply the bottom eq by 5 to cancel out the x when u add the equations

OpenStudy (anonymous):

x-35y+35-2y=-2

OpenStudy (anonymous):

could the answer be (1,-7)

OpenStudy (anonymous):

no doesn't fit in

OpenStudy (anonymous):

AHHHHHH!!!!!!!!

OpenStudy (anonymous):

wait maybe i can help @ganeshie8 @mathmate

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@AravindG @YanaSidlinskiy

OpenStudy (anonymous):

are you finding people who might know?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ppl are nice here they might help

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

iight cool, its not like any of my teachers

OpenStudy (anonymous):

my teachers never wanted to help

ganeshie8 (ganeshie8):

-5(7y-7)-2y=-2 -35y + 35 - 2y = -2 -37y = -37 y = ?

OpenStudy (anonymous):

-1

ganeshie8 (ganeshie8):

think again

ganeshie8 (ganeshie8):

-37y = -37 divide 37 both sides -y = -1 multiply -1 both sides y = 1

OpenStudy (anonymous):

oh, duh

ganeshie8 (ganeshie8):

substitute this value in any on of the given equations and solve x

OpenStudy (anonymous):

thanks @ganeshie8

OpenStudy (anonymous):

yeah, thank you @ganeshie8

ganeshie8 (ganeshie8):

np :) maybe substituting in second equation is slightly easier : x - 7y = -7 x -7(1) = -7 x - 7 = 7 x = 0

ganeshie8 (ganeshie8):

so the solution of given system is ` (0, 1) `

ganeshie8 (ganeshie8):

Notice that we always write the x coordinate first.

OpenStudy (anonymous):

ok now i get it too

OpenStudy (anonymous):

yeah i knew the (x,y). some how i had (1,-7)

OpenStudy (anonymous):

what about this one? Solve this system of equations. Write your answer as an ordered pair. y + 6 = x 2x + 2y = 0

OpenStudy (anonymous):

hey @ganeshie8 , how did you get -37